OpenStudy (anonymous):

3x + 6x-45 =180 Solve for x

OpenStudy (anonymous):

first step is to bring the 180 to the left hand side and collect like terms.

OpenStudy (anonymous):

do that first.

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

9x-45

OpenStudy (anonymous):

what about -45-180?

OpenStudy (anonymous):

3x + 6x -45 = 180 9x -45 -180 = 0 9x -225=0 now you try the rest.

OpenStudy (anonymous):

Your confusing me even more. I dont get what im supposed to do

OpenStudy (anonymous):

tell me what u are confused about please? and are u even reading what I say?

OpenStudy (anonymous):

Yes i am.

OpenStudy (anonymous):

in the question it clearly says "Solve for x"

OpenStudy (anonymous):

you want to find x, and in this case x = .......... (the dots are to be replaced with a value that you will find by solving.

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

No;(

OpenStudy (anonymous):

oh god.

OpenStudy (anonymous):

3x + 6x -45 = 180, do you agree? that is just the equation

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

now collect like terms. 3x + 6x = 9x u did that at first but u forgot about the -45 and -180. 3x + 6x -45=180 so lets start over by bringing the 180 to the left hand side. 3x + 6x - 45 -180=0 9x -225 = 0 9x = 225 x = 225/9 x = 25

OpenStudy (anonymous):

I listed every step there and it's the most simplest form so if u don't understand just ask me what u are confused about otherwise I cannot help u further because that's all there is to it.

OpenStudy (anonymous):

25 isn't an option

OpenStudy (anonymous):

well obviously u made a mistake typing the equation/question or the options given by ur textbook or whoever it may be are wrong.

OpenStudy (anonymous):

please tell me the options.

OpenStudy (anonymous):

supplementary means sum = 180 so, 3x + 6x-45 =180 now solve for 'x'

OpenStudy (anonymous):

Thats the equation i came up with but heres the problem.

OpenStudy (anonymous):

sorry, the whole diagram and question cannot be seen and it's blurry.

OpenStudy (anonymous):

if that's an equation u came up with, then most likely it's wrong then, please show me the full question.

OpenStudy (anonymous):

I didn't come up with it.

OpenStudy (anonymous):

 8. Given that line f is parallel to line g, find the value of x. The diagram is not to scale.